OpenStudy (anonymous):

Am I doing this wrong? What is the solution of log2x − 6 256 = 4? I'm trying to find the value of x but when I plug in x nothing equals out to 4.

OpenStudy (anonymous):

\[\log(2x-6)*256=4\]

OpenStudy (anonymous):

x = −5 x = −4 x = 4 x = 5 Here are my choices.

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

2x-6 is the base, right ?

OpenStudy (anonymous):

Correct

hartnn (hartnn):

\(\huge \log_{(2x-6)}256=4\)

hartnn (hartnn):

(2x-6)^4 = 256

hartnn (hartnn):

4^4 = 256

hartnn (hartnn):

which gives 2x-6 = 4

hartnn (hartnn):

now get x

OpenStudy (anonymous):

Okay.. But wouldn't be easier to just plug in the x values into the equation or would it not work?

OpenStudy (anonymous):

And x would equal 5

hartnn (hartnn):

sometimes plugging in is easier, sometimes its not. if you know log rules, then this is much easier to actually solve rather than trying out all choices

OpenStudy (anonymous):

Okay, I appreciate the help!

hartnn (hartnn):

now verify plug in 5

OpenStudy (anonymous):

its 5